OpenStudy (anonymous):

Find all solutions of the equation. (Enter each answer in the form θ + 2πk, 0 ≤ θ < 2π. If there are any unused answer boxes, enter NONE in the last boxes.) 2cos^2(x)-cos(x)-1=0

OpenStudy (anonymous):

You can factor this expression.

OpenStudy (anonymous):

i factored it to get (2cos(x) + 1)(cos(x) - 1) = 0

OpenStudy (anonymous):

\[(2\cos(x)+1)(\cos(x)-1)=0\]

OpenStudy (anonymous):

now set each factor equal to zero and solve for x

OpenStudy (anonymous):

Cos(x)= -1/2 and Cos(x)= 1

OpenStudy (anonymous):

yes, but now you need to express all possible angles for which those ratios are correct

OpenStudy (anonymous):

2pi/3, 4pi/3 and 2pi?

OpenStudy (anonymous):

plus 2pik on all of them

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you!